OpenStudy (anonymous):

solve equation 1/5 + x/8 = 6/7

OpenStudy (anonymous):

Subtract both sides by 1/5 and multiply by 8.

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

It's an algebra question. Follow the steps Ahaanomegas has provided you with to find the answer.

OpenStudy (anonymous):

what sides do i subtract?

OpenStudy (anonymous):

Do each one to both sides! Your first step is: \[\frac{ 1 }{ 5 }+ \frac{ x }{ 8} - \frac{ 1 }{ 5 } = \frac{ 6 }{ 7 } -\frac{ 1 }{ 5 }\] Subtracting 1/5 from both sides. Then you simplify that so that you just have x/8 on the left, and then you multiply by 8 on both sides so that you just have x on the left.

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

No problem. Ask again if you have more questions.

OpenStudy (anonymous):

k

hero (hero):

1/5 + x/8 = 6/7 x/8 = 6/7 - 1/5 x/8 = 30/35 - 7/35 x/8 = 23/35 Cross multiply: 35x = 8*23 x = 8*23/35 x = 184/35

hero (hero):

Any questions?

OpenStudy (anonymous):

Wow, thank u soo much